OpenStudy (anonymous):

Solve √(9x+81)=x+5 The solution is x=

OpenStudy (anonymous):

is it -8,7 and do I put parenthesis around them

OpenStudy (anonymous):

Square both sides:\[9x+81 = (x+5)^{2}\]\[9x+81 = x^{2}+10x+25\]Get everything on one side:\[0 = x^{2}+x-56\]Factor:\[0 = (x-7)(x+8)\]So x is 7 or -8

OpenStudy (mayliz):

x=7

OpenStudy (freckles):

7(9)+81=63+81=144 7+5=12 7 works ------ 9(-8)+81=9 -8+5=-3 -8 doesn't work ------- :) Great work!